ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 256, 2023

Aharoni, R. and H. Guo Rainbow cycles for families of matchings  1
Alon, N. Spanning trees with few non-leaves  9
Alon, N. and R. Holzman Near-sunflowers and focal families  21
Bárány, I., G. Kalai and A. Pór Universal sequences of lines in ℝ^{d}  35

Becker, O., A. Lubotzky and J. Mosheiff Testability in group theory

61

Beniamini, G. and N. Nisan Bipartite perfect matching as a real polynomial

91

Chudnovsky, M. and G. Kalai Attempting perfect hypergraphs  133
Even-Zohar, C. and J. Hass Random colorings in manifolds  153

Farber, M. and T. Nowik The homology of random simplicial complexes in the multi-parameter upper model

213

Friedgut, E. and Y. Iluz Hyper-regular graphs and high dimensional expanders

233

Friedman, J. and D. Puder A note on the trace method for random regular graphs

269

Goldreich, O. and A. Wigderson Good permutation codes based on the shuffle-exchange network

283

Guo, H. See R. Aharoni Guterman, A., R. Meshulam and I. Spiridonov Maximal gener-

alized rank in graphical matrix spaces

297

Hass, J. See C. Even-Zohar Holzman, R. See N. Alon

Hrubeš. P. and A. Yehudayoff Shadows of Newton polytopes  311

Iluz, Y. See E. Friedgut Kalai, G. See I. Bárány Kalai, G. See M. Chudnovsky Keller, C. and M. A. Perles No Krasnosel’skiĭ number for general

sets

345

Kwan, M., A. Sah, M. Sawhney and M. Simkin Substructures in Latin squares

363

Lee, J. R. Spectral dimension, Euclidean embeddings, and the metric growth exponent

417

Lee, T. and A. Shraibman Around the log-rank conjecture  441

Lew, A., E. Nevo, Y. Peled and O. E. Raz On the d-dimensional algebraic connectivity of graphs

479

Li, Y., Y. Qiao, A. Wigderson, Y. Wigderson and C. Zhang Connections between graphs and matrix spaces

513

Lubotzky, A. See O. Becker Meshulam, R., I. Newman and Y. Rabinovich Large simple d-

cycles in simplicial complexes

581

Meshulam, R. See A. Guterman Mosheiff, J. See O. Becker Naor, A., A. Sah, M. Sawhney and Y. Zhao Cayley graphs that

have a quantum ergodic eigenbasis

599

Nevo, E. See A. Lew Newman, I. See R. Meshulam Nisan, N. See G. Beniamini Nowik, T. See M. Farber Peled, Y. See A. Lew Perles, M. A. See C. Keller

Pinchasi, R. On the odd area of the unit disc  619

Pór, A. See I. Bárány Puder, D. See J. Friedman Qiao, Y. See Y. Li Rabinovich, Y. See R. Meshulam Raz, O. E. See A. Lew Sah, A. See A. Naor Sah, A. See M. Kwan Samorodnitsky, A. One more proof of the first linear programming

bound for binary codes and two conjectures

639

Sawhney, M. See A. Naor Sawhney, M. See M. Kwan Shraibman, A. See T. Lee Simkin, M. See M. Kwan Spiridonov, I. See A. Guterman Wagner, U. and P. Wild Coboundary expansion, equivariant over-

lap, and crossing numbers of simplicial complexes

675

Wigderson, A. See O. Goldreich Wigderson, A. See Y. Li Wigderson, Y. See Y. Li Wild, P. See U. Wagner Yehudayoff, A. See P. Hrubeš Zhang, C. See Y. Li Zhao, Y. See A. Naor